Radison Enterprises sells a product for $108 per unit. The variable cost is $56 per unit, while fixed costs are $1,297,920.

Determine (a) the break-even point in sales units and (b) the break-even point if the selling price were increased to $116 per unit.

a. Break-even point in sales units units
b. Break-even point if the selling price were increased to $116 per unit units
Answer #1
a. Break-even point in units =             24,960 units
b. Break-even point in units =             21,632 units
Workings:
a. Unit selling Price = $               108
Unit variable cost = $                 56
Unit contribution margin = $                 52
Break-even point in units = Fixed costs / Unit contribution margin
= $1297920 / $52
=             24,960
b. Unit selling Price = $               116
Unit variable cost = $                 56
Unit contribution margin = $                 60
Break-even point in units = Fixed costs / Unit contribution margin
= $1297920 / $60
=             21,632
Answer #2

SOLUTION :


a.



Contribution margin per unit (UCM) = SP - UVC = 108 - 56 = 52 ($)


BEP (units) = FC / UCM = 1297920 / 52 = 24960 units (ANSWER)





b.


If SP = $116 :


Contribution margin per unit (UCM) = SP - UVC = 116 - 56 = 60 ($)


BEP (units) = FC / UCM = 1297920 / 60 = 21632 units (ANSWER)

answered by: Tulsiram Garg